# Learning UAV-based path planning for efficient localization of objects using prior knowledge

### v1

@@ 2.3 Policy learning
where $Q(s,a)$ is the learned action-value function. Q-learning is a popular model-free method that learns the action-value function by iteratively optimizing a Q-table using the immediate reward $r_{t}$ and the discounted future reward $\gamma\cdot\max\limits_{a\in A}Q(s_{t+1},a)$, where $\gamma\in[0,1]$ is the discount factor determining the emphasis on the future reward (Sutton and Barto, 2018). Q-Learning is, however, unsuitable for learning high-dimensional state-spaces because the size of the Q-table grows exponentially with the number of possible states and actions. To overcome this problem, we used a Deep Q-Network (DQN) introduced by Mnih et al. (2015). DQN uses a neural network to approximate the action-value function. The architecture of this network is described in section 2.3.1. Section 2.3.2 describes the experience replay buffer which is used to provide training samples for the network.The training procedure is described in section 2.3.3.
@@ 2.3.1 Network architecture
Figure 4 shows the used network architecture for the Q-network. The network gets the current state $s_{t}$ as input and predicts the Q-value for every action $a\in A$. The network consists of a feature extractor and a fully connected network. The feature extractor consist of two parallel convolution blocks to ensure unique feature extraction from both local and global map. The output of both convolution blocks is flattened and concatenated with the movement budget to combine the extracted features of the local and global map with the movement budget. Four fully-connected layers convert these features into four outputs which correspond to the action values for each possible action $a\in A$. This architecture is based on the architecture used in Theile et al. (2021).
Figure 4: Network architecture for the DQN using the global and local map and the movement budget indicating the input size, the number of kernels and their size, the size of the flatten layer and the size of the fully connected layers. The number of trainable parameters equals 2,544,548.
@@ 2.3.2 Experience Replay Buffer
To create training data for the DQN, we used an experience replay buffer (Mnih et al., 2015). The replay buffer is a circular buffer of size $n_{\textrm{buffer}}$ containing state transition vectors, $[s_{t},a_{t},r_{t},s_{t+1}]$. From this buffer, mini-batches of size $n_{\textrm{batch}}$ are sampled for training the Q-network. The buffer is filled by an agent that operates in the simulation following a probability-based training policy that takes random actions from action space $A$ using probability vector $p$:
$\forall a\in A,p(s,a)=\frac{\exp(Q(s,a)/\tau)}{\sum\limits_{\forall a_{i}\in A}\exp(Q(s,a_{i})/\tau)},$ (2)
where $\tau\in(0,\infty)$ is the temperature parameter. A high value for $\tau$ gives all actions equal probability and a low value results in a greedy training policy (Equation 1). This probability-based training policy has an advantage over a standard $\epsilon$-greedy policy because $\tau$ is independent of the number of training steps, whereas $\epsilon$ usually requires a schedule to decrease with the number of training steps. The training policy is implemented in a similar way as Theile et al. (2020).
2.3.3 Training procedure
Figure 5 shows the training procedure for the DQN. During training, two identical neural networks are used: a policy network $Q$ and a target network $Q_{\textrm{target}}$. The weights of the target network are replaced by the weights of the policy network every $n_{\textrm{update}}$ training steps. This is done to stabilise the learning by reducing the correlation between the action values $Q(s,a)$ and the target value (immediate reward + discounted future rewards) (Mnih et al., 2015). The optimal action-value function is approximated by minimizing the smooth L1 loss (Girshick, 2015) between the target, $y=r_{t}+\gamma\max\limits_{a_{i}\in A}(Q_{\textrm{target}}(s_{t+1},a_{i}))$, and the predicted value, $\hat{y}=Q(s_{t},a_{t})$, using $\beta=1$. Optimizing the weights of Q is done using an Adam optimizer with learning rate $\alpha$.
@@ 2.3.3 Training procedure
Figure 5: Training procedure for the Deep Q-Network.
Table 2: Hyper-parameters for training the Deep Q-Network.
@@
$\tau$ 0.005 Temperature parameter
$\gamma$ 0.95 Discount factor
@@ 2.4 Experiments
In this section, we present the experiments to validate the learned search policy. We evaluate the percentage of objects found and flight path length. These values are compared to a traditional row-by-row coverage flight path, planned by Fields2Cover (Mier et al., 2023), using the size of the FoV, $F$, as path width without overlap. Because we expected the learned path planner to outperform a row-by-row flight path specifically for non-uniformly distribution objects, we evaluated the impact of these object distributions (section 2.4.1). To test the robustness of the RL-learned path planner, we evaluated the effect of detection errors (section 2.4.2) and prior knowledge quality (section 2.4.3). Lastly, the effect of different stopping criteria was evaluated to mark the end of the search (section 2.4.4).
2.4.1 Experiment 1: Impact of object distributions
To study the influence of the object distribution on the path length and percentage of found objects, three object distributions were defined. Distribution ’strong’ was a strong clustered distribution and used the default field parameters from Table 1. Distribution ’medium’ was a medium clustered distribution with $\mathcal{N}_{\textrm{dist}}(\mu,\sigma)=\mathcal{N}(4,1)$ and $\mathcal{N}(\mu_{i},\Sigma_{i})$ has a random mean $\mu_{i}$ and a covariance $\Sigma_{i}$ uniformly sampled from set $\{\Sigma_{1},\Sigma_{2},\Sigma_{3},\Sigma_{4}\}$ where $\Sigma_{1}=[\begin{smallmatrix}10&16\\ 16&40\end{smallmatrix}]$, $\Sigma_{2}=[\begin{smallmatrix}40&0\\ 0&10\end{smallmatrix}]$, $\Sigma_{3}=[\begin{smallmatrix}30&12\\ 12&12\end{smallmatrix}]$, and $\Sigma_{4}=[\begin{smallmatrix}15&4\\ 4&20\end{smallmatrix}]$. Distribution ’uniform’ was a uniform distribution with each object having a uniform random non-overlapping coordinate in the field. For all distributions, the number of objects in the field was drawn from the same normal distribution $\mathcal{N_{\textrm{obj}}}(\mu,\sigma)$. An example of each distribution is shown in Figure 6. A different policy was trained and evaluated using 1000 evaluation episodes for the three distributions.
Figure 6: Example of a field with (a) a strong distribution, (b) a medium distribution and (c) a uniform distribution of objects.
@@ 2.4.2 Experiment 2: Influence of detection errors
To assess the influence of errors of the simulated detection network on the policy, we defined five detection error levels, ranging from a very high number of errors to no errors. The parameters for each level of detection errors are given in Table 3. The default simulation environment from table 1 is equivalent to the moderate level in table 3. Figure 7 shows an example of the simulated output of the detection network. For each level of detection errors, a policy was trained and evaluated using 1000 evaluation episodes.
@@ 2.4.3 Experiment 3: Influence of prior knowledge quality
We assessed two aspects of quality in the prior knowledge: the uncertainty related to resolution and the inaccuracy in the prior knowledge. When the resolution of the prior knowledge is low, only sparse information about the location of the objects is available. This could, for example, be information from other sources with a lower spatial resolution such as satellite images or images taken from a higher altitude. Another aspect of prior knowledge is the inaccuracy. The higher the inaccuracy, the less reliable the prior knowledge becomes. This could for example be due to sensor noise or detection errors. We defined five levels of prior knowledge quality, ranging from no prior knowledge (level none) to perfect prior knowledge (level perfect), expressing an increasing quality. Table 4 defines these levels. The default simulation environment from table 1 is equivalent to the moderate level in table 4. To keep the same number of trainable parameters in the DQN, we kept the same number of cells in the prior knowledge map for all quality levels by using nearest neighbor upsampling to resize the prior knowledge map to $48\times 48$. Figure 8 shows an example of the prior knowledge map for each level. For each quality level, a different policy was trained and evaluated using 1000 evaluation episodes.
@@
Figure 8: Example of the prior knowledge of world (a) for quality level none (no prior knowledge) (b), low (c), moderate (d), high (e), and perfect (f). Level perfect has high-resolution knowledge about the locations of the objects, while level low has limited knowledge about the locations of objects. For visibility, a colormap is applied.
@@ 2.4.4 Experiment 4: Effect of different stopping criteria
In a real-world application, the number of objects might be unknown which makes the termination of the simulation when all objects are found infeasible. Therefore, we evaluated the effect of different stopping criteria, specifically:
@@
    Stop searching when there are no new objects detected for a certain number of flight actions. To avoid constantly resetting this counter due to FP detections, at least 2 objects need to be detected. Three thresholds were tested, 15, 25 and 50 steps.
@@
These stopping criteria were compared with the default stopping criterion which stops searching when all objects are found. We compared both the percentage of found objects and the path length. For each stopping criterion, a separate policy was trained and evaluated on 1000 evaluation episodes.
@@ 3 Results
3.1 Experiment 1: Impact of object distributions
Figure 9 shows the relation between flight path length and the percentage of found objects for different distributions of objects. The more uniformly distributed the objects were, the more linear the relation between path length and number of found objects. For the strong and medium distributions, the learned policy outperformed the baseline row-by-row flight path, having found more than 80% of the objects in 73 and 94 steps on average respectively, compared to 209 steps for the baseline row-by-row flight path. However the learned policy often had troubles to find all objects before the battery was empty.
Figure 9: Effect of a strong, a medium and a uniform weed distribution on both the number of found objects and the path length for the policy learned by DQN, and the baseline row-by-row flight path. The lines show the mean over 1000 episodes.
Figure 10 shows examples of a single flight path for the strong, medium and uniform distributions, respectively. It can be seen that the drone flew in quite straight lines till around 80% of the objects were found. After 80% the drone started wandering around till the battery was empty and crashed, because the drone missed some objects and environment did only terminate when all objects were found.
Figure 10: Single flight paths of the RL agent (a-c) and the baseline row-by-row flight path (d-f) for strong (a,d), medium (b,e) and uniform (c,f) distribution of objects. The detected objects are indicated with black dots and the not detected objects with gray dots. The red line indicates the flight path till 80% of the objects are found, the blue line the complete flight path till all objects were found or the battery was empty.
@@ 3.2 Experiment 2: Influence of detection errors
Figure 11 shows the relation between the different levels of detection errors and the percentage of found objects and flight path length. All the levels outperformed the baseline row-by-row flight path. The learned policy is robust and only starts to drop performance for a very high number of detection errors. All levels outperform the baseline by having found 80% of the objects in 70 - 98 steps compared to 206 steps for the baseline row-by-row flight path.
Figure 11: Effect of the different levels of detection errors on the percentage of found objects for the policy learned by DQN, and the baseline row-by-row flight path. The lines show the mean over 1000 episodes.
Figure 12 shows some flight paths for the different levels of detection errors. For a very high number of detection errors (Figure 12a), there are many false positive detections visible. When having a perfect detection network (Figure 12e), the agent missed some objects, however, this is just an example of a single flight path. On average, this level of detection errors performed comparable to the levels low-high (Figure 11).
Figure 12: Single flight paths of the RL agent for detection error levels very low (a), low (b), moderate (c), high (d) and perfect (e) and the baseline row-by-row flight path (f). The detected objects are indicated with black dots and the not detected objects with gray dots. The red line indicates the flight path till 80% of the objects are found, the blue line the complete flight path till all objects were found or the battery was empty.

### v2

@@ 2.3 Policy learning
where $Q(s,a)$ is the learned action-value function. Q-learning is a popular model-free method that learns the action-value function by iteratively optimizing a Q-table using the immediate reward $r_{t}$ and the discounted future reward $\gamma\cdot\max\limits_{a\in A}Q(s_{t+1},a)$, where $\gamma\in[0,1]$ is the discount factor determining the emphasis on the future reward (Sutton and Barto, 2018). Q-Learning is, however, unsuitable for learning high-dimensional state-spaces because the size of the Q-table grows exponentially with the number of possible states and actions. To overcome this problem, we used a Deep Q-Network (DQN) introduced by Mnih et al. (2015). DQN uses a neural network to approximate the action-value function. Although more RL algorithms exist, we use DQN because of the discrete action space and its sample efficiency (relative to other RL algorithms). However, the RL algorithm can easily be changed since the simulation environment uses the standard OpenAI Gym API.
Figure 4 shows the training procedure for the DQN. During training, two identical neural networks are used: a policy network $Q$ and a target network $Q_{\textrm{target}}$. Both networks have an equal architecture, which is described in section 2.3.1. During training, the parameters from the target network, $\theta_{\textrm{target}}$, are updated by the policy network parameters, $\theta$, using a soft update:
$\theta_{\textrm{target}}\leftarrow(1-\tau)\theta_{\textrm{target}}+\tau\theta,$ (2)
where $\tau\in[0,1]$ defines the update rate. This is done to stabilize the learning by reducing the correlation between the action values $Q(s,a)$ and the target value (immediate reward + discounted future rewards) (Mnih et al., 2015). Training data for the DQN is created using a replay buffer, which is explained in section 2.3.2. The optimization of the policy network is described in section 2.3.3.
Figure 4: Training procedure for the Deep Q-Network.
@@ 2.3.1 Network architecture
Figure 5 shows the network architecture used for the Q-network. This architecture is based on the architecture used in Theile et al. (2021). The network gets the current state $s_{t}$ as input and predicts the Q-value for every action $a\in A$. The network consists of a feature extractor and a fully connected network. The feature extractor consists of two parallel convolution blocks to ensure unique feature extraction from both the local and the global map. The output of both convolution blocks is flattened and concatenated with the movement budget to combine the extracted features of the local and global map with the movement budget. Four fully-connected layers convert these features into four outputs that correspond to the action values for each possible action $a\in A$. Compared to Theile et al. (2021), the input global map was larger (after down-sampling, see Section 2.2), resulting in a larger flattened layer and an increased number of trainable parameters.
Figure 5: Network architecture for the DQN using the global and local map and the movement budget indicating the input size, the number of kernels and their size, the size of the flatten layer, and the size of the fully connected layers. The number of trainable parameters equals 2,544,548.
@@ 2.3.2 Experience Replay Buffer
To create training data for the DQN, we used an experience replay buffer (Mnih et al., 2015). The replay buffer is a circular buffer of size $n_{\textrm{buffer}}$ containing state transition vectors, $[s_{t},a_{t},r_{t},s_{t+1}]$. From this buffer, mini-batches of size $n_{\textrm{batch}}$ are sampled for training the Q-network. The buffer is filled by an agent that operates in the simulation following a probability-based training policy that takes random actions from the action space $A$ using the probability vector $p$:
$\forall a\in A,p(s,a)=\frac{\exp(Q(s,a)/\lambda)}{\sum\limits_{\forall a_{i}\in A}\exp(Q(s,a_{i})/\lambda)},$ (3)
where $\lambda\in(0,\infty)$ is the temperature parameter. A high value for $\lambda$ gives all actions equal probability, and a low value results in a greedy training policy (Equation 1). This probability-based training policy has an advantage over a standard $\epsilon$-greedy policy because $\lambda$ is independent of the number of training steps, whereas $\epsilon$ usually requires a schedule to decrease with the number of training steps. The training policy is equal to Theile et al. (2020).
2.3.3 Optimizing policy network
The optimal action-value function is approximated by minimizing the smooth L1 loss (Girshick, 2015) between the target, $y=r_{t}+\gamma\max\limits_{a_{i}\in A}(Q_{\textrm{target}}(s_{t+1},a_{i}))$, and the predicted value, $\hat{y}=Q(s_{t},a_{t})$, using $\beta=1$. Optimizing the weights of Q is done using an Adam optimizer with learning rate $\alpha$.
@@ 2.3.3 Optimizing policy network
Table 2: Hyperparameters for training the Deep Q-Network.
@@
$\lambda$ 0.1 Temperature parameter
@@ 2.4 Experiments
In this section, we present the experiments to validate the learned search policy. We evaluate the percentage of weeds found and the flight-path length. These values are compared to a traditional row-by-row coverage flight path, planned by Fields2Cover (Mier et al., 2023), using the size of the FoV, $F$, as path width without overlap. Because we expected the learned path planner to outperform a row-by-row flight path specifically for non-uniformly distributed weeds, we evaluated the impact of these weed distributions (section 2.4.1). To test the robustness of the RL-learned path planner, we evaluated the effect of detection errors (section 2.4.2) and prior knowledge quality (section 2.4.3). We also investigated the effect of different stopping criteria to mark the end of the search (section 2.4.4). Lastly, the transferability of the simulation-trained RL policy was evaluated on real-world image data (section 2.4.5).
2.4.1 Experiment 1: Impact of weed distributions
To study the influence of the weed distribution on the path length and percentage of found weeds, three weed distributions were defined. Distribution ’strong’ was a strong clustered distribution and used the default field parameters from Table 1. Distribution ’medium’ was a medium clustered distribution with $\mathcal{N}_{\textrm{dist}}(\mu,\sigma)=\mathcal{N}(4,1)$ and $\mathcal{N}(\mu_{i},\Sigma_{i})$ has a random mean $\mu_{i}$ and a covariance $\Sigma_{i}$ uniformly sampled from set $\{\Sigma_{1},\Sigma_{2},\Sigma_{3},\Sigma_{4}\}$ where $\Sigma_{1}=[\begin{smallmatrix}10&16\\ 16&40\end{smallmatrix}]$, $\Sigma_{2}=[\begin{smallmatrix}40&0\\ 0&10\end{smallmatrix}]$, $\Sigma_{3}=[\begin{smallmatrix}30&12\\ 12&12\end{smallmatrix}]$, and $\Sigma_{4}=[\begin{smallmatrix}15&4\\ 4&20\end{smallmatrix}]$. Distribution ’uniform’ was a uniform distribution with each weed having a uniform random non-overlapping coordinate in the field. For all distributions, the number of weeds in the field was drawn from the same normal distribution $\mathcal{N_{\textrm{obj}}}(\mu,\sigma)$. An example of each distribution is shown in Figure 6. All other parameters were kept at their default value as defined in Table 1 and 2. A different policy was trained and evaluated using 1000 evaluation episodes for the three distributions.
Figure 6: Example of a field with (a) a strong distribution, (b) a medium distribution, and (c) a uniform distribution of weeds.
@@ 2.4.2 Experiment 2: Influence of detection errors
To assess the influence of errors of the simulated detection network on the policy, we defined five detection error levels, ranging from a very high number of errors to no errors. The parameters for each level of detection errors are given in Table 3. The default simulation environment from Table 1 is equivalent to the moderate level in Table 3. Figure 7 shows an example of the simulated output of the detection network. All other parameters were kept at their default value as defined in Table 1 and 2. For each level of detection errors, a policy was trained and evaluated using 1000 evaluation episodes.
@@ 2.4.3 Experiment 3: Influence of prior knowledge quality
We assessed two aspects of quality in the prior knowledge: the uncertainty related to resolution and the inaccuracy in the prior knowledge. When the resolution of the prior knowledge is low, only sparse information about the location of the weeds is available. This could, for example, be information from other sources with a lower spatial resolution, such as satellite images or images taken from a higher altitude. Another aspect of prior knowledge is the inaccuracy. The higher the inaccuracy, the less reliable the prior knowledge becomes. This could, for example, be due to sensor noise or detection errors. We defined five levels of prior knowledge quality, ranging from no prior knowledge (level none) to perfect prior knowledge (level perfect), expressing an increasing quality. Table 4 defines these levels. The default simulation environment from Table 1 is equivalent to the moderate level in Table 4. To keep the same number of trainable parameters in the DQN, we kept the same number of cells in the prior knowledge map for all quality levels by using nearest neighbor upsampling to resize the prior knowledge map to $48\times 48$. Figure 8 shows an example of the prior knowledge map for each level. All other parameters were kept at their default value as defined in Table 1 and 2. For each quality level, a different policy was trained and evaluated using 1000 evaluation episodes.
@@
Figure 8: Example of the prior knowledge of world (a) for quality level none (no prior knowledge) (b), low (c), moderate (d), high (e), and perfect (f). Level perfect has high-resolution knowledge about the locations of the weeds, while level low has limited knowledge about the locations of weeds. For visibility, a colormap is applied.
@@ 2.4.4 Experiment 4: Effect of different stopping criteria
In a real-world application, the number of weeds might be unknown, which makes the termination of the simulation when all weeds are found infeasible. Therefore, we evaluated the effect of different stopping criteria, specifically:
@@
    Stop searching when there are no new weeds detected for a certain number of flight actions. To avoid constantly resetting this counter due to FP detections, at least 2 weeds need to be detected. Three thresholds were tested: 15, 25, and 50 steps.
@@
These stopping criteria were compared with the default stopping criterion, which stops searching when all weeds are found. We compared both the percentage of found weeds and the path length. For each stopping criterion, a separate policy was trained and evaluated on 1000 evaluation episodes.
2.4.5 Experiment 5: Transferability of simulation-learned policy to a real-world application
To test the transferability of the learned policy to a real-world scenario, the planner was tested on real-world image data without further training. To do so, we evaluated its performance on the four real-world datasets presented in van Essen et al. (2025). Each dataset consists of a high-resolution orthomosaic of a grass field containing artificial plants distributed in clusters in the field, as a proxy for a weed detection application. Images were captured at four different dates in spring and summer using a DJI M300 drone with a Zenmuse P1 RGB camera, each date having a different weed distribution. The orthomosaics were made using Agisoft Metashape (Agisoft, 2023) using the high-quality settings. Compared to van Essen et al. (2025), we only use a part of the orthomosiacs to match the squared input format of the RL agent.
Figure 9 illustrates the generation of the local weed map using real-world orthomosaic data. To align with the state representation of the RL agent, we rasterize the fields in a raster of $M\times M$ grid cells with a grid-cell size of 1x1 m. At each time step, a camera image was generated by mapping the UAV position to real-world coordinates, cropping an area of $F\times F$ m around this position from the orthomosaic, and resizing the crop to an image of 2048×2048 pixels using linear interpolation. Weeds were detected in the camera images using YOLOv8-nano (Jocher et al., 2023). All detected weeds with a confidence score larger than 0.5 were mapped to the local map coordinates to create the local weed map. The detection network was trained for 250 epochs on the training dataset used in (van Essen et al., 2025), which consisted of real-world drone images from the grass field with artificial plants taken from 12m, 24m, and 32m altitude. In total, 1618 images with 2000 annotations were available. Compared to (van Essen et al., 2025), both classes were merged into a single ’plant’ class.
Figure 9: Applying the simulation-trained Reinforcement Learning policy to real-world image data by generating an image from the orthomosaic (with the raster shown in red) corresponding to the drone’s field of view (blue rectangle), and converting it into a local weed map using an object detection network.
To create the global map, prior knowledge was generated using a high-altitude row-by-row flight path with a field-of-view of 24x24 grid cells. In these images, weeds were detected using the trained detection network. All detected weeds with a confidence score larger than 0.05 were mapped to the global map coordinates.
For the experiment, we used the trained RL policy with learned landing action from experiment 4. The results of this policy on the four datasets were compared to the baseline row-by-row flight path. We compared both the percentage of found weeds and the flight-path length.
@@ 3 Results
3.1 Experiment 1: Impact of weed distributions
Figure 10 shows the relation between flight-path length and the percentage of found weeds for different distributions of weeds. The more uniformly distributed the weeds were, the more linear the relation between path length and the number of found weeds. For the strong and medium distributions, the learned policy outperformed the baseline row-by-row flight path, having found more than 80% of the weeds in 73 and 94 steps on average, respectively, compared to 209 steps for the baseline row-by-row flight path. However, the learned policy often had trouble finding all weeds before the battery was empty.
Figure 10: Effect of a strong, a medium, and a uniform weed distribution on both the number of found weeds and the path length for the policy learned by DQN, and the baseline row-by-row flight path. The lines show the mean over 1000 episodes.
Table 5 shows the percentage of found weeds at 100, 200, and 300 flight steps for the policy learned by DQN and the baseline row-by-row flight path. Up to 200 flight steps, the DQN policy significantly found more weeds than the baseline. However, in contrast to the baseline, it did not find all weeds.
Table 5: Percentage of found weeds at 100, 200, and 300 flight steps for a strong, a medium, and a uniform weed distribution for the policy learned by DQN, and the baseline row-by-row flight path. The values show the mean ± the standard deviation. Values indicated with a ’*’ have a significant ($\alpha=0.001)$ higher percentage of found weeds than the baseline row-by-row flight path (Welch’s t-test).
Method 100 steps 200 steps 300 steps
row-by-row 0.37±0.08 0.76±0.07 1.00±0.00
dqn-strong 0.88±0.14* 0.95±0.08* 0.97±0.06
dqn-medium 0.82±0.13* 0.93±0.08* 0.95±0.05
dqn-uniform 0.55±0.06* 0.81±0.07* 0.89±0.05
Figure 11 shows examples of a single flight path for the strong, medium, and uniform distributions, respectively. It can be seen that the drone flew in quite straight lines till around 80% of the weeds were found. After 80% the drone started wandering around till the battery was empty and crashed, because the drone missed some weeds, and the environment only terminated when all weeds were found.
Figure 11: Single flight paths of the RL agent (a-c) and the baseline row-by-row flight path (d-f) for strong (a,d), medium (b,e), and uniform (c,f) distribution of weeds. The detected weeds are indicated with black dots and the undetected weeds with gray dots. The red line indicates the flight path till 80% of the weeds are found, the blue line the complete flight path till all weeds were found or the battery was empty.
@@ 3.2 Experiment 2: Influence of detection errors
Figure 12 shows the relation between the different levels of detection errors and the percentage of found weeds and flight-path length. All the levels outperformed the baseline row-by-row flight path. The learned policy is robust and only starts to drop performance for a very high number of detection errors. All levels outperform the baseline by having found 80% of the weeds in 70–98 steps compared to 206 steps for the baseline row-by-row flight path.
Figure 12: Effect of the different levels of detection errors on the percentage of found weeds for the policy learned by DQN, and the baseline row-by-row flight path. The lines show the mean over 1000 episodes.
Table 6: Percentage of found weeds at 100, 200, and 300 flight steps for different levels of detection errors for the policy learned by DQN, and the baseline row-by-row flight path. The values show the mean ± the standard deviation. Values indicated with a ’*’ have a significant ($\alpha=0.001)$ higher percentage of found weeds than the baseline row-by-row flight path (Welch’s t-test).
Method 100 steps 200 steps 300 steps
row-by-row 0.37±0.30 0.76±0.25 1.00±0.00
dqn-very high 0.80±0.17* 0.92±0.08* 0.94±0.06
dqn-high 0.90±0.11* 0.96±0.05* 0.97±0.04
dqn-moderate 0.90±0.12* 0.96±0.05* 0.97±0.04
dqn-low 0.90±0.13* 0.96±0.05* 0.97±0.04
dqn-perfect 0.90±0.13* 0.95±0.07* 0.96±0.06
Table 6 shows the percentage of found weeds at the different detection error levels corresponding to 100, 200, and 300 flight steps for the DQN policy and the baseline row-by-row flight path. The differences between the levels were small, only the very-high level of detection errors performed slightly lower, although it still found significantly more weeds than the baseline until 200 flight steps. Up to 200 flight steps, the DQN policy had found significantly more weeds than the baseline.
Figure 13 shows some flight paths for the different levels of detection errors. For a very high number of detection errors (Figure 13a), there are many false positive detections visible. When having a perfect detection network (Figure 13e), the agent missed some weeds, however, this is just an example of a single flight path. On average, this level of detection errors performed comparable to the levels low-high (Figure 12).
Figure 13: Single flight paths of the RL agent for detection error levels very low (a), low (b), moderate (c), high (d), and perfect (e) and the baseline row-by-row flight path (f). The detected weeds are indicated with black dots and the undetected weeds with gray dots. The red line indicates the flight path till 80% of the weeds are found, the blue line the complete flight path till all weeds were found or the battery was empty.